how many molecules of carbon dioxide are contained in one litre of the air if the volume content of carbon dioxide is 0.03% at ntp?
 

Grade:11

1 Answers

Sakshi
8 years ago
At NTP, 1 L of air has CO2 = 0.03*1L= 0.03L
Applying the ideal gas equation, p*V= n*R*T
Putting the values NTP as 293.15 K, and 1 atm, one can calculate the value of n
multiplying that value of n with the avegadro number gives the number of molecules of CO2 present